# EXGnet: a single-lead explainable-AI guided multiresolution network with train-only quantitative features for trustworthy ECG arrhythmia classification

### v1

@@ 2.2 Proposed Network Architecture
The proposed framework’s architectural representation is depicted in figure 2. This framework includes XAI guidance, features used exclusively during training, and a novel base architecture including multiresolution block. The architecture of the framework, along with the motivation behind each design choice, is explained in detail in the following sections.
@@ 2.2.1 Base network with multiresolution feature extraction
Our proposed network takes a 10-second ECG signal as input, capturing both short-term and long-term features inherent in the signal. Figure 2 shows the detailed block diagram of our proposed method. Unlike most existing approaches that focus exclusively on either short-term or long-term features, or employ separate CNN networks to extract and integrate both [21], our design takes a better approach. We introduce a residual multiresolution block that utilizes convolutions with varying kernel sizes and depth-wise convolutions to capture multiresolution features without the need for separate CNNs. To address the computational overhead associated with larger kernels, we draw inspiration from Navid et.al. and Showrav et.al. [37, 38], implementing a parallel convolutional structure. Here, we use three 16$\times$16 kernel convolutions to get the advantage of both 16$\times$16, 31$\times$31, and 46$\times$46 convolutions, thus reducing the computation cost. This design achieves multiresolution feature extraction with reduced computational complexity.
The residual multiresolution block also includes depth-wise convolution, enabling the network to learn features from individual channels. Similar to the work of Hannun et.al. [11], we incorporate shortcut connections to facilitate efficient optimization. Our base network comprises 16 residual multiresolution blocks, with every alternate block subsampling its inputs by a factor of two. Corresponding Max Pooling layers within the blocks also perform subsampling on alternate layers. Batch normalization is applied before each ReLU activation layer, and a Dropout layer is used to mitigate overfitting. The number of channels per multiresolution block begins with 32 and doubles after every four blocks, ending with 256 in the final block.
@@ 2.2.2 XAI guidance
In this paper, we propose a novel approach that incorporates XAI-guided training. Specifically, we use Grad-CAM [39] within the network to generate CAMs from the output of the final multiresolution block. In our model, the CAM consists of 20 values, each corresponding to a sequential block of 250 samples from the original 5000-sample ECG signal. A CAM value close to 1 indicates that the corresponding region contains features critical to the final prediction. Previous research has shown that many severe heart conditions are associated with significant variability in the RR interval [40]. Therefore, we initially focus on RR interval variability in the ECG signal. To guide the network, we apply a 1D mask of 20 samples long generated by considering the RR intervals provided by the NeuroKit 0.2 algorithm, which is designed to capture heart rate variability. By guiding the network with these masks, the model learns to concentrate on clinically relevant regions of the signal, thereby emulating the decision-making process of a cardiologist.
@@ 2.2.3 Train-only quantitative features integration
Machine learning approaches to ECG arrhythmia classification rely on the extraction of various statistical, morphological, and handcrafted features from the ECG signal, followed by the utilization of classifiers such as Random Forest, Gradient Boosting, and Naive Bayes. However, these methods frequently fall short of providing optimal results. To improve performance, some studies have integrated these quantitative features with deep learning networks, yielding better outcomes. Despite these advances, the complexity of calculating numerous features poses a significant challenge for portable ECG devices, where computational efficiency is crucial.
To overcome this limitation, we utilized train-only quantitative features, demonstrating that better results can be achieved also when these features are excluded during testing. In the testing phase, only the raw ECG signal is required, eliminating the computational burden of feature extraction, and making the system more suitable for portable ECG devices. We employ 17 distinct quantitative features, which pass through a tiny, separate network. This network consists of three dense layers, with the final layer’s size corresponding to the number of arrhythmia classes and utilizing Softmax activation. The output of the second dense layer is concatenated with the output of the GAP layer of the base network. The combined output then passes through another dense layer with Softmax activation for final prediction. This strategic use of train-only quantitative features allows the network to learn additional features through a reciprocal loss function, enhancing the overall performance of the model.
@@ 2.3 Loss functions
where, $N$ is the number of samples, $C$ is the number of classes, and $\hat{y}_{i,c}$ is the predicted probability where the ground truth probability is ${y}_{i,c}$. Besides, to guide the network with XAI, we use Grad-CAM. To guide the CAM values we utilize Normalized Cross Correlation (NCC) loss function, which helps the network to learn making the CAM values orientation as per the ground truth. The NCC loss can be expressed as
@@
3 Results
3.1 Experimental setup
3.1.1 Implementation details
3.1.2 Evaluation metrics
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (6)
$\text{Sensitivity}=\frac{TP}{TP+FN}$ (8)
$\text{Specificity}=\frac{TN}{TN+FP}$ (9)
@@ 3.1.2 Evaluation metrics
3.2 Performance comparison
@@ 4.1 Ablation study
We conducted a comprehensive ablation study to quantitatively assess the impact of each new component in our proposed method. Table 4 outlines the incremental performance improvements achieved through various configurations: (a) Baseline, (b) Baseline with multiresolution feature extraction, (c) Baseline + multiresolution feature extraction + XAI guidance via Grad-CAM, (d) Baseline + multiresolution feature extraction + quantitative features integration, (e) Baseline + multiresolution feature extraction + XAI guidance + quantitative features integration (results from the base network EXGnet), and (f) Baseline + multiresolution feature extraction + XAI guidance + quantitative features integration (results from the combined network EXGnet-L). We present the ablation study on the both Chapman and Ningbo datasets, with average five fold results evaluated using Accuracy, F1-score, Sensitivity, and Specificity, as shown in Table 4.
@@ 4.2 Evaluation of trustworthiness
A similar scenario is observed in sample (e), where the left CAM indicates that the network did not focus on the most critical features of the ECG signal. In contrast, with XAI guidance, the model learns to concentrate on the relevant areas that cardiologists typically consider when making decisions. Thus, Figure 3 clearly illustrates the qualitative impact of XAI guidance and underscores its importance in achieving reliable predictions.
@@ 4.3 Limitations and future works
Our proposed network shows reliable and promising results when compared with state-of-the-art approaches. It is not without limitations, though. First, the number of quantitative features trained could be expanded to allow the model to learn more statistical characteristics. Second, we relied solely on Grad-CAM for XAI guidance due to its simplicity. While effective, other advanced CAM-generating techniques might enhance the network’s ability to focus on critical features of the ECG signal.
@@ 5 Conclusion
In this paper, we proposed EXGnet, a single-lead XAI guided network designed for trustworthy ECG arrhythmia classification. We introduced a multiresolution feature extraction method that reduces computational demands while maintaining performance. By employing Grad-CAM based XAI guidance alongside train-only quantitative features, we aimed to enhance the model’s performance and trustworthiness. Our findings demonstrate that XAI guidance significantly improves the network’s ability to detect arrhythmias from true abnormal regions in the ECG signal. While the machine learning features contribute positively, they are not necessary during testing primarily; however, their inclusion during testing leads to a notable performance increase. According to the experimental results, EXGnet outperforms recent state-of-the-art networks across all metrics. Overall, our investigation confirms that this technique is not only effective but also reliable, making it well-suited for practical applications in portable single-lead ECG devices and beyond.
@@ Acknowledgements
Author contributions
Tushar Talukder Showrav: Original draft preparation, Methodology, Software, Validation, Investigation, Visualization. Soyabul Islam Lincoln: Original draft preparation, Software, Validation, Investigation. Md. Kamrul Hasan: Data curation, Conceptualization, Supervision, Writing – review & editing.
Funding
The authors received no specific funding for this study.
Availability of data and materials
We used two public datasets in this work. Both of these datasets are jointly available online in a Physionet official link (https://physionet.org/content/ecg-arrhythmia/1.0.0/).
Ethics approval and consent to participate
Not applicable.
Consent for publication
Not applicable.
Competing Interests
The authors declare that there are no competing interests.

### v2

@@ 2.2 Proposed Network Architecture
The proposed framework’s architectural representation is depicted in figure 3. This framework includes XAI guidance, features used exclusively during training, and an innovative base architecture including multiresolution block. The architecture of the framework, along with the motivation behind each design choice, is explained in detail in the following sections.
@@ 2.2.1 Base network with multiresolution feature extraction
Our proposed network takes a 10-second ECG signal as input, capturing both short-term and long-term features inherent in the signal. Figure 3 shows the detailed block diagram of the proposed method. Unlike most existing approaches that focus exclusively on either short-term or long-term features, or employ separate CNN networks to extract and integrate both [28], the proposed design takes a better approach. We introduce a residual multiresolution block that utilizes convolutions with varying kernel sizes and depth-wise convolutions to capture multiresolution features without the need for separate CNNs. To address the computational overhead associated with larger kernels, we draw inspiration from Navid et.al. and Showrav et.al. [36, 37], implementing a parallel convolutional structure. Here, three 16$\times$16 kernel convolutions are used to get the advantage of both 16$\times$16, 31$\times$31, and 46$\times$46 convolutions, thus reducing the computation cost. This design achieves multiresolution feature extraction with reduced computational complexity. This process can be expressed as
[…]
where $\psi^{k}$ denotes a convolutional layer with kernel size $k$, and $\parallel$ denotes concatenation. The residual multiresolution block also includes depth-wise convolution, enabling the network to learn features from individual channels. Similar to the work of Hannun et.al. [38], shortcut connections are incorporated to facilitate efficient optimization. The base network comprises 16 residual multiresolution blocks, with every alternate block subsampling its inputs by a factor of two. Corresponding Max Pooling layers within the blocks also perform subsampling on alternate layers. Batch normalization is applied before each ReLU activation layer, and a Dropout layer is used to mitigate overfitting. The number of channels per multiresolution block begins with 32 and doubles after every four blocks, ending with 256 in the final block.
@@ 2.2.2 XAI guidance
In this paper, we introduce an innovative approach that incorporates XAI guidance directly into the training process. Our method integrates a CAM generation mechanism, inspired by Grad-CAM [39], within the network itself. Specifically, CAMs are produced from the output of the final multiresolution feature block, allowing the model to highlight signal regions that contribute most significantly to its predictions. The complete algorithm for CAM generation is outlined in Algorithm 1. In our model, the CAM consists of 20 values, each corresponding to a sequential block of 250 samples from the original 5000-sample ECG signal. A CAM value close to 1 indicates that the corresponding region contains features critical to the final prediction. Previous research has shown that many severe heart conditions are associated with significant variability in the RR interval [40]. Therefore, we initially focus on RR interval variability in the ECG signal. To guide the network, we apply a 1D mask of 20 samples long generated by considering the RR intervals provided by the NeuroKit2 algorithm, which is designed to capture heart rate variability. By guiding the network with these masks, the model learns to concentrate on clinically relevant regions of the signal, thereby emulating the decision-making process of a cardiologist.
Algorithm 1 Network-integrated CAM generation
1:Input: Model $M$, Input Signal $X$, Target Feature Layer $L$
2:Output: Class Activation Map (CAM) $H$
3:
4:procedure GenerateCAM($M,X,L$)
5:     $A,Y_{scores}\leftarrow\text{ForwardPass}(M,X,L)$
6:     $c\leftarrow\arg\max(Y_{scores})$
7:     $y^{c}\leftarrow Y_{scores}[c]$
8:     $G\leftarrow\frac{\partial y^{c}}{\partial A}$triangleright Compute gradients w.r.t. feature map
9:     $P\leftarrow A\odot G$triangleright Element-wise product
10:     $H_{raw}\leftarrow\frac{1}{N_{channels}}\sum_{k}P_{k}$triangleright Average across channels
11:     $H\leftarrow\frac{H_{raw}-\min(H_{raw})}{\max(H_{raw})-\min(H_{raw})}$triangleright Normalize to [0, 1]
12:     return $H$
13:end procedure
@@ 2.2.3 Train-only quantitative features integration
Machine learning approaches to ECG arrhythmia classification rely on the extraction of various statistical, morphological, and handcrafted features from the ECG signal, followed by the utilization of classifiers such as Random Forest, Gradient Boosting, and Naive Bayes. However, these methods frequently fall short of providing optimal results. To improve performance, some studies have integrated these quantitative features with deep learning networks, yielding better outcomes. Despite these advances, the complexity of calculating numerous features poses a significant challenge for edge ECG devices, where computational efficiency is crucial.
To address this limitation, we introduce the use of train-only quantitative features, demonstrating that superior performance can still be achieved even when these features are excluded during inference. During testing, only the raw ECG signal is required, thereby eliminating the computational overhead of feature extraction and enhancing the system’s suitability for edge ECG devices. Specifically, we extract 17 distinct quantitative features, which are processed through a compact auxiliary network (Feature-based network) composed of three fully connected (dense) layers. The final layer of this network corresponds to the number of arrhythmia classes and employs a Softmax activation function. The output from the second dense layer is concatenated with the output of the GAP layer from the base ECG network. This fused representation is subsequently passed through an additional dense layer with Softmax activation to generate the final prediction.
The overall process can be expressed as:
$y_{j}=J\left(\overline{G(x)},\overline{F(x_{f})}\right),$ (4)
where $G(\cdot)$ denotes the base network operating on the raw ECG input $x$, $F(\cdot)$ represents the feature network processing the quantitative feature vector $x_{f}$, and $J(\cdot)$ denotes the joint network responsible for final classification. Here, $\overline{G(x)}$ and $\overline{F(x_{f})}$ represent intermediate outputs of the base and auxiliary networks, respectively, which are fused for the final prediction. Although the primary emphasis of our approach lies in the base network, the quantitative features are not available during testing. However, the strategic incorporation of these train-only features during training enables the base network to learn enriched representations. This is achieved through a reciprocal loss mechanism, ultimately improving the overall classification performance.
@@ 2.3 Loss functions
where, $N$ is the number of samples, $C$ is the number of classes, and $\hat{y}_{i,c}$ is the predicted probability where the ground truth probability is ${y}_{i,c}$. Besides, to guide the network with XAI, we use a network-integrated CAM generation technique. To guide the CAM values we utilize NCC loss function, which helps the network to learn making the CAM values orientation as per the ground truth. The NCC loss can be expressed as
@@
3 Results
3.1 Experimental setup
3.1.1 Implementation details
3.1.2 Evaluation metrics
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (8)
$\text{Sensitivity}=\frac{TP}{TP+FN}$ (10)
$\text{Specificity}=\frac{TN}{TN+FP}$ (11)
Figure 4: Visualization of the aggregated confusion matrix of the proposed network on both datasets.
@@ 3.2 Performance comparison
To further observe the classification performance, Figure 4 shows the aggregated confusion matrices for the EXGnet model on both the Chapman and Ningbo datasets. For the Chapman dataset (Figure 4a), the confusion matrix shows excellent performance, with high values concentrated along the main diagonal, indicating a high rate of correct classifications for all five arrhythmia types. Misclassifications are minimal, demonstrating the model’s robustness. The confusion matrix for the Ningbo dataset (Figure 4b) also shows strong results, though it reflects the more challenging nature of this data. The most prominent confusion is observed between Sinus Rhythm and Sinus Irregularity, with 154 Sinus Irregularity samples misclassified as Sinus Rhythm and 112 Sinus Rhythm signals misclassified as Sinus Irregularity. Despite these challenges, the model maintains compelling accuracy across most classes, reinforcing the overall superior performance reported in Table 2.
Figure 5: t-SNE visualization of the feature maps extracted from the immediate last layer in different ablation study cases for Fold 1 of the Chapman dataset. The cases are: (a) Baseline, (b) a + Multiresolution feature extraction, (c) b + XAI guidance, (d) b + Quantitative features, (e) b + c + d (EXGnet), and (f) b + c + d (EXGnet-L).
@@ 4.1 Ablation study
We conducted a comprehensive ablation study to quantitatively assess the impact of each new component in our proposed method. Table 3 outlines the incremental performance improvements achieved through various configurations: (a) Baseline, (b) Baseline with multiresolution feature extraction, (c) Baseline + multiresolution feature extraction + XAI guidance, (d) Baseline + multiresolution feature extraction + quantitative features integration, (e) Baseline + multiresolution feature extraction + XAI guidance + quantitative features integration (results from the base network EXGnet), and (f) Baseline + multiresolution feature extraction + XAI guidance + quantitative features integration (results from the combined network EXGnet-L). We present the ablation study on the both Chapman and Ningbo datasets, with average five fold results evaluated using Accuracy, F1-score, Sensitivity, and Specificity, as shown in Table 3.
@@
To provide a qualitative counterpart to our quantitative results, we visualized the high-dimensional feature space using t-SNE. Figure 5 illustrates the resulting feature distributions for each model configuration on fold 1 of the Chapman dataset. Figure 5(a) shows the feature distribution for the baseline model, where the clusters for different rhythm classes, especially Sinus Rhythm (red) have significant overlap with other classes. With the addition of the multiresolution feature extraction block in Figure 5(b), the clusters begin to show better separation and become more compact, visually confirming the performance gains reported in Table 3. The addition of XAI guidance (Figure 5(c)) and quantitative features (Figure 5(d)) further improves these clusters, making them tighter and reducing overlapping cases.
A clear improvement is visible in Figure 5(e), which corresponds to our proposed EXGnet model. Here, the feature clusters are much more distinct and well-defined, showing that the network learned features that can better distinguish between the classes. Finally, Figure 5(f) shows the feature space for the complete EXGnet-L model. The clusters are extremely tight and clearly separated, with large margins between different classes. This visual evidence strongly supports the numerical results, showing that each proposed component helps in creating a more separable feature representation, which leads to the superior classification performance of the final model.
Table 4: Quantitative analysis of the interpretability of the proposed network.
Dataset Metric Without XAI With XAI
Chapman NCC $0.0181\pm 0.0056$ $0.6844\pm 0.0125$
L1 $0.7523\pm 0.0174$ $0.2154\pm 0.0076$
L2 $0.8008\pm 0.0110$ $0.3472\pm 0.0115$
Ningbo NCC $0.0813\pm 0.0072$ $0.9192\pm 0.0454$
L1 $0.5662\pm 0.0307$ $0.1411\pm 0.0106$
L2 $0.6368\pm 0.0355$ $0.2915\pm 0.0211$
Figure 7: Potential anomalous cases identified from the Ningbo dataset. (a) Real label is Sinus Rhythm, but predicted as Sinus Irregularity. (b) Real label is Sinus Irregularity, but predicted as Sinus Rhythm.
Figure 8: t-SNE visualization of the feature maps extracted from the immediate last layer for Fold 5 of the Ningbo dataset.
@@ 4.2 Evaluation of trustworthiness
A similar scenario is observed in sample (e), where the left CAM indicates that the network did not focus on the most critical features of the ECG signal. In contrast, with XAI guidance, the model learns to concentrate on the relevant areas that cardiologists typically consider when making decisions. Thus, Figure 6 clearly illustrates the qualitative impact of XAI guidance and underscores its importance in achieving reliable predictions. To quantitatively assess the trustworthiness of the proposed technique, we introduced an innovative evaluation approach. Specifically, the CAMs generated by Grad-CAM are compared with ground truth CAMs using three metrics: NCC, L1 distance, and L2 distance. As shown in Table 4, the inclusion of XAI guidance during training significantly improves the alignment between predicted and true CAMs across both datasets. Specifically, the NCC scores increase substantially, indicating a higher similarity, while the L1 and L2 errors decrease, reflecting better localization accuracy and reduced deviation. These results demonstrate the effectiveness of the proposed XAI-guided model in producing more trustworthy and interpretable explanations.
4.3 Confusing cases identified in the Ningbo dataset
Figure 4(b) reveals that a significant portion of signals from Ningbo dataset are confused between Sinus Rhythm and Sinus Irregularity. To further explore this observation, we examined the t-SNE visualization in Figure 8, based on fold 5 of the Ningbo dataset. This plot reflects how the network separates different classes in the high-dimensional feature space. Notably, there is a considerable overlap between the Sinus Rhythm and Sinus Irregularity classes, suggesting that the network perceives similar features in some instances of both categories. While it is true that these cases are inherently challenging for the model, further inspection revealed the presence of certain anomalies that could be contributing to this confusion. We observed that a noticeable number of signals labeled as Sinus Rhythm exhibit features commonly associated with Sinus Irregularity, and vice versa. Representative examples from the validation set of fold 5 are shown in Figure 7.
Figure 7(a) presents two signals with ground truth labels of Sinus Rhythm. However, upon closer observation, subtle regions of irregular rhythm are visible, which likely led the network to predict them as Sinus Irregularity. In contrast, Figure 7(b) displays two signals labeled as Sinus Irregularity that were predicted as Sinus Rhythm. Interestingly, these signals show only minimal rhythm irregularities, even less pronounced than those in Figure 7(a). Such borderline cases challenge the model’s ability to learn class-distinctive features effectively.
It is important to note that an ideal Sinus Irregular Rhythm should display a noticeable variation in heart rhythm. Another type of anomaly found in the dataset involves incorrect labeling between Sinus Bradycardia and Sinus Rhythm, as well as between Sinus Tachycardia and Sinus Rhythm. For example, the second signal in Figure 7(a) appears to have a heart rate of approximately 120 beats per minute, yet it is labeled as Sinus Rhythm instead of Sinus Tachycardia. These mislabeled or ambiguous cases can significantly mislead the network during training. Although limitations exist within the model itself, achieving higher performance and reliability ultimately depends on the accuracy and consistency of the dataset labels.
@@ 4.4 Limitations and future works
Our proposed network shows reliable and promising results when compared with state-of-the-art approaches. It is not without limitations, though. First, the number of quantitative features trained could be expanded to allow the model to learn more statistical characteristics. Second, we relied solely on a Grad-CAM inspired CAM generation technique for XAI guidance due to its simplicity. While effective, other advanced CAM-generating techniques might enhance the network’s ability to focus on critical features of the ECG signal.
@@ 5 Conclusion
In this work, we introduced EXGnet, a reliable and trustworthy deep learning network developed for single-lead ECG arrhythmia classification, with a strong focus on enhancing interpretability through XAI guidance during training. By supervising the learning process with ground truth CAMs generated through an innovative heart rate variability based technique, the model is guided to focus on clinically meaningful regions of the ECG signal. This guidance not only improves classification performance but also significantly enhances the trustworthiness and transparency of the model’s decision-making process. To enrich the model’s learning with multi-domain knowledge, we incorporated quantitative features during training. These train-only features provide the model with complementary information beyond what is learned from raw data. While these features are not required during testing, their optional inclusion at inference leads to notable performance improvements. Furthermore, we designed a multiresolution feature extraction strategy to capture relevant patterns at different temporal scales. This design choice enables the model to maintain high performance while keeping computational demands low, making it well suited for real-time applications in edge environments. Our extensive experiments demonstrate that EXGnet surpasses recent state-of-the-art models across all key metrics. These results highlight the potential of EXGnet for practical deployment in portable and edge-based single-lead ECG monitoring systems, including smartphones and wearable healthcare devices.
@@ Acknowledgements
CRediT authorship contribution statement
Tushar Talukder Showrav: Original draft preparation, Writing – review & editing, Conceptualization, Methodology, Software, Validation, Investigation, Visualization. Soyabul Islam Lincoln: Original draft preparation, Software, Validation, Investigation. Md. Kamrul Hasan: Conceptualization, Supervision, Writing – review & editing.
Data Availability Statement
No new data were generated or analyzed during this study.
Conflict of Interest
The authors state that they have no conflicts of interest.